Circular No. 6700

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1997cx IN NGC 3057 Michael Schwartz, Cottage Grove, OR, reports his discovery of a su-

pernova (mag ∼ 15) located ∼ 20′′ south of the center of NGC 3057 (α = 10^{h}05^{m}.7, δ = +80^{o}17′, equinox 2000.0) on a CCD image taken with the 0.35-m automated Tenagra Observatory supernova patrol telescope on July 12.33 UT (with a confirming image taken an hour later).

S. Jha, P. Garnavich, P. Challis, and R. Kirshner, Harvard-Smithson-ian Center for Astrophysics, report that a spectrogram taken by P. Berlind at the Whipple Observatory 1.5-m telescope on July 14.2 UT indicates that SN 1997cx is a type-II supernova past maximum, showing P-Cyg Balmer lines, He I (587.5 nm), and Fe II. The expansion velocity measured from the Hα profile is 7000 km/s. Images taken concurrently by C. W. Hergenrother at the 1.2-m telescope yield magnitude R = 15.4 and the following position: α = 10^{h}05^{m}37^{s}.43, δ = +80^{o}17′27′′.6 (equinox 2000.0).

SUPERNOVA 1997cw IN NGC 105 Garnavich, Jha, Kirshner, and Challis also report that spectra of SN

1997cw were obtained by P. Berlind with the Whipple Observatory 1.5-m telescope on July 14.5 UT, confirming that this is a type-Ia supernova. However, the depth of the broad Na I absorption and Fe II lines shortward of 550 nm suggest that the supernova is at least two weeks past maximum brightness, significantly older than reported on IAUC 6699. The Si II 615.0nm absorption is indeed weak, and the spectrum is similar to that of SN 1991T taken 15 days past maximum (Phillips et al. 1992, A.J. 103, 1632). Images taken near the time of the spectra, by Hergenrother at the 1.2-m telescope, yield magnitude V = 16.4.

Corrigendum. On IAUC 6699, line 14, for 3.7 nm) read 0.37 nm)

BL LACERTAE M. Maesano, E. Massaro, and R. Nesci, Astronomical Institute, Uni-

versity ‘La Sapienza’, Rome, report that BL Lac is still in outburst (cf. IAUC 6693), exhibiting rapid and strong variations (ΔV = 1.54 over 1 day). Johnson B and V and Cousins R and I magnitudes: July 12.93 UT, 15.42, 14.52, 13.88, 13.10; July 13.96, 13.84, 12.98, 12.36, 11.65.

1997 July 14  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green